OpenStudy (anonymous):

D. Simplify: square root 30y^3 divided by square root 5xy

OpenStudy (anonymous):

\[\frac{30y^3}{5xy}=\frac{6y^2}{x}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):